# Quasi-boolean groups

### v2

@@ 4. Profinite principal $G$-bundles
Let $A$ be a topological space and let $m\colon A\to B$ be a map. The quotient topology on $B$ with respect to $m$ is defined the following way: a subset $U\subseteq B$ is open if and only if $m^{-1}(U)\subseteq A$ is open. Now let $A$ be a topological space equipped with a left action of a group $H$. Let $H\backslash A$ denote the quotient of $A$ with respect to the left action of $H$ equipped with the quotient topology.
@@ 1. $(a)$
    Let $A$ be a topological space and let $m\colon A\to B$ be a map. Let $B$ be equipped with the quotient topology with respect to $m$. Then a map $h\colon B\to C$ of topological spaces is continuous if and only if the composition $h\circ m$ is continuous.
@@ 2. $(b)$
    Let $A$ be a topological space equipped with a left action of a group $H$. Then the quotient map $A\to H\backslash A$ is open, that is, it maps open sets to open sets.
@@ 3. $(c)$
    Let $A,B$ be two topological spaces both equipped with a left action of a group $H$. Then the product topology on $H\backslash A\times H\backslash B$ is the quotient topology with respect to the quotient map
    $A\times B\longrightarrow(H\times H)\backslash(A\times B)=H\backslash A\times H\backslash B.$
@@
is continuous with respect to the product topologies, the pre-image of any open subset is open. Therefore we only need to show that if $(q_{A}\times q_{B})^{-1}(U)\subseteq A\times B$ for a subset $U\subseteq H\backslash A\times H\backslash B$ is open, then $U$ is open in the product topology. In this case, $(q_{A}\times q_{B})^{-1}(U)$ is the union of sets of the form $V\times W$, where $V\subseteq A$ and $W\subseteq B$ are open. By part $(b)$ the set $(q_{A}\times q_{B})(V\times W)=q_{A}(V)\times q_{B}(W)$ is open in the product topology. Since $U$ is the union of such subsets, it is open, and $(c)$ follows. ∎
@@
Let $X$ be a profinite space and let $\mathcal{U}$ be an open covering of $X$. Then there is a finite open covering $\mathcal{V}$ of $X$ consisting of pairwise disjoint open and closed subsets which is subordinate to $\mathcal{U}$.
@@
Because $X$ is profinite, its open and closed subsets form a subbasis of its topology. Therefore there is an open covering $\mathcal{W}$ of $X$ consisting of open and closed subsets which is subordinate to $\mathcal{U}$. Since $X$ is compact, we may assume without the loss of generality that $\mathcal{W}$ is finite. Let $R$ be the subring of $\mathbf{B}(X)$ generated by the characteristic functions of the elements of $\mathcal{U}$. Since $R$ is finitely generated, it is finite by Corollary 2.11.
Let $\mathcal{V}$ be the collection of the characteristic functions of the atoms of $R$. We claim that $\mathcal{V}$ satisfies the required properties. Because $R$ is finite, the set $\mathcal{V}$ is also finite. Since different atoms of $R$ are orthogonal by part $(iii)$ of Lemma 2.15, their support is pair-wise disjoint by part $(i)$ of Lemma 2.15. Let $e_{1},e_{2},\ldots,e_{n}$ be the atoms of $R$. Then $1=e_{1}+\ldots+e_{n}$, so the union of their supports is $X$. Also the elements of $\mathcal{V}$ are open, since they are the supports of elements of $\mathbf{B}(X)$.
Now let $e$ be an atom of $R$ and pick an $x$ in its support $V$. Since $\mathcal{U}$ is a covering, there is a $U\in\mathcal{U}$ such that $x\in U$. The characteristic function $f$ of $U$ is in $R$, so it is the sum of different atoms of $R$. Therefore, $ef$ is either 00 or $e$ by the orthogonality of atoms. In the first case, the intersection of the supports $U$ and $V$ is empty, but both contain $x$, a contradiction. Hence $ef=e$, so $U$ contains $V$, and hence $\mathcal{V}$ is subordinate to $\mathcal{U}$. ∎
Recall that a section of a continuous map of topological spaces $f\colon Y\to X$ is a continuous map $s\colon X\to Y$ such that $f\circ g$ is the identity map of $X$.
@@
By assumption there is an open cover $\mathcal{U}$ of $X$ such that for every $x\in X$ there is a $U\in\mathcal{U}$ such that $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$ has a section. By Lemma 4.3 there is a finite open covering $\mathcal{V}$ of $X$ consisting of pairwise disjoint open and closed subsets which is subordinate to $\mathcal{U}$. Let $V\in\mathcal{V}$ and pick a $U\in\mathcal{U}$ which contains $V$. By assumption there is a section of $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$; its restriction to $V$ is a section $s_{V}\colon V\to f^{-1}(V)$ of $f|_{f^{-1}(V)}\colon f^{-1}(V)\to V$. Since the elements of $\mathcal{V}$ of $X$ form an open and pairwise disjoint covering of $X$, the union $\bigcup_{V\in\mathcal{V}}s_{V}$ of these sections is a section of $f\colon Y\to X$. ∎
@@
Let $G$ be a compact group, and let $X$ be a profinite space equipped with a free continuous group action $g\colon G\times X\to X$. Then both $G$ and $G\backslash X$ are profinite.
@@
For the sake of simple notation, in the sequel we will denote any left action of any group on any set by multiplication on the left, if this does not lead to confusion. Since the action $g$ is continuous and free, for every $x\in X$ the map $\gamma\mapsto\gamma x$ from $G$ onto the $G$-orbit of $x$ is continuous and injective. As $G$ is compact and $X$ is Hausdorff, we get that the image of this map is closed, and the map itself is a homeomorphism from $G$ onto its image. As $G$ is homeomorphic to a closed subspace of a profinite space, it is also profinite by Theorem 3.2.
Let $q\colon X\to G\backslash X$ denote the quotient map. Since $q$ is surjective, continuous, and $X$ is compact, we get that $G\backslash X$ is compact, too. Let $x,y\in G\backslash X$ be two arbitrary different points. The pre-images $C=q^{-1}(x)$ and $D=q^{-1}(y)$ are disjoint $G$-orbits, and they are also closed by the above. Therefore by Lemma 3.6 there exist disjoint open and closed subsets $U,V\subset X$ such that $U$ contains $C$ and $V$ contains $D$.
Set $U^{\prime}=\bigcup_{\gamma\in G}\gamma U$. Since each $\gamma\in G$ acts as a homeomorphism, we get that $U^{\prime}$ is the union of open subsets, so it is open. It is also the continuous image of $G\times U$. Both $G,U$ are compact since the latter is closed in a profinite space. Hence their product $G\times U$ is also compact. Thus, $U^{\prime}$ is closed, since $Y$ is Hausdorff. Also $U^{\prime}\cap D$ is empty; if it were not then $\gamma U\cap D\neq\emptyset$ for some $\gamma\in G$, but then $U\cap D=\gamma^{-1}(\gamma U\cap D)\neq\emptyset$ using that $D$ is $G$-invariant. This is a contradiction. A similar argument shows that $V^{\prime}=\bigcup_{\gamma\in N}\gamma U$ is open and closed, and $V^{\prime}\cap C$ is empty.
Therefore $W=U^{\prime}\cap V^{\prime}$ is open and closed, and $W$ is disjoint from $C\cup D$. Therefore its complement $U^{\prime\prime}=U^{\prime}-W$ in $U^{\prime}$ is also open and closed, and contains $C$ since $U^{\prime}$ does. Similarly $V^{\prime\prime}=V^{\prime}-W$ is also open and closed, and contains $D$. Moreover, both $U^{\prime}$ and $V^{\prime}$ are $G$-invariant, so the same holds for $W$, and hence for $U^{\prime\prime}$ and $V^{\prime\prime}$. Because these sets are disjoint, their images $q(U^{\prime\prime})$ and $q(V^{\prime\prime})$ are also disjoint. They are also open by part $(b)$ of Lemma 4.2, and since $x,y$ are in $q(U^{\prime\prime})$ and $q(V^{\prime\prime})$, respectively, and were arbitrary, we get that $G\backslash X$ is at least Hausdorff.
Going back to the situation above, since $U^{\prime\prime}$ and $V^{\prime\prime}$ are closed in a profinite space, they are compact. Hence their images $q(U^{\prime\prime})$ and $q(V^{\prime\prime})$ are closed, since $G\backslash X$ is Hausdorff. Therefore they are a pair of disjoint open and closed neighbourhoods of $x$ and $y$; since the latter were arbitrary, we get that $G\backslash X$ is even totally separated, so it is profinite by Theorem 3.2. ∎
@@
Let $X$ be a topological space and let $G$ be a compact group. A profinite principal $G$-bundle over $X$ is a continuous map $f\colon Y\to X$ equipped with a free continuous group action $g\colon G\times Y\to Y$ such that $Y$ is profinite, the map $f$ is surjective and the pre-image of each $x\in X$ is a $G$-orbit.
@@
$(i)$ In the situation of the theorem above both $G$ and $X$ are profinite. This is immediate for $G$ from Proposition 4.5. Moreover there is a unique bijection $\iota\colon G\backslash Y\to X$ such that $f=\iota\circ q$, where $q\colon Y\to G\backslash Y$ denotes the quotient map. Since $f$ is continuous, we get that $\iota$ is a continuous bijection by part $(a)$ of Lemma 4.2. Since $G\backslash Y$ is compact by Proposition 4.5 and $X$ is Hausdorff, we get that $\iota$ is a homeomorphism. Moreover $G\backslash Y$ is actually profinite by Proposition 4.5, so the same holds for $X$, too.
$(ii)$ The reader might wonder if it is true that for every continuous, surjective map $f\colon X\to Y$ of profinite spaces have a section. It turns out that those $Y$ which have this property for every such $f$ have a name: extremally disconnected spaces. They can be characterised by the following property: every set of open and closed subsets of $Y$ has a supremum with respect to inclusion (see the Folk Theorem of [3] on page 485). Some spaces, such as the Čech–Stone compactification of discrete spaces have this property, but there are many profinite spaces which do not.
$(iii)$ The theorem above was already stated by Morel, see the remark after Lemma 4 of [7] on page 359. However, no proof was given, just a remark that the strategy of the proof of Proposition 1 of [12] on page 4 works. This is what we will do, but for the sake of the reader we will give a detailed argument.
@@
By Proposition 4.4 it will be sufficient to show that every $x\in X$ has an open neighbourhood $U$ such that $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$ has a section. Since $f^{-1}(x)$ is homeomorphic to $G$, it is non-empty, so there is a $y\in Y$ such that $f(y)=x$. Because $G$ is finite, and every point of $Y$ is closed, as $Y$ is Hausdorff by Theorem 3.2, there exist pairwise disjoint open and closed subsets $U_{\gamma}\subset X$ for all $\gamma\in G$ such that $U_{\gamma}$ contains $\gamma y$ for each $\gamma$ by Lemma 3.6.
Set $V=\bigcap_{\gamma\in G}\gamma^{-1}U_{\gamma}$. It is the intersection of finitely many open and closed subsets, so it is also open and closed. For every $\gamma\in G$ we have $\gamma V\subseteq\gamma(\gamma^{-1}U_{\gamma})=U_{\gamma}$, so $V\cap\gamma V=\emptyset$ for every $\gamma\neq 1$ in $G$. Therefore the restriction of $f$ to $V$ is injective. Since $V$ is closed, it is compact, while its image $f(V)$ is Hausdorff, so $f|_{V}$ has a continuous inverse $f(V)\to V$. Since $y=\gamma^{-1}(\gamma y)\in\gamma^{-1}U_{\gamma}$, we get that $y\in\gamma^{-1}U_{\gamma}$ for every $\gamma\in G$, so $y\in V$, and hence $x\in f(V)$.
Therefore it will be sufficient to show that $f(V)$ is open in $X$. Set $Z=\bigcup_{\gamma\in G}\gamma V$; it is clearly $G$-invariant, and since it is the union of open subsets, it is also open. Therefore its complement $W=Y-Z$ in $Y$ is $G$-invariant and closed. Moreover $f(V)=f(Z)$, as $f$ is $G$-equivariant, and $f$ is surjective, so the complement of $f(V)$ in $X$ is $f(W)$. Since $W$ is closed, it is compact, and $X$ is Hausdorff, so $f(W)$ is closed. Therefore its complement $f(V)$ is open. ∎
Definition 4.10.
Let $N$ be a closed, normal subgroup of $G$ and let $G_{N}=N\backslash G$ denote the quotient: it is a profinite group. Let $r_{N}\colon G\to G_{N}$ be the quotient map, which is a continuous group homomorphism. There is a unique map $f_{N}\colon N\backslash Y\to X$ such that $f\colon Y\to X$ is the composition of the quotient map $q_{N}\colon Y\to N\backslash Y$ and $f_{N}$. There is a unique group action $G\times N\backslash Y\to N\backslash Y$ which makes $f_{N}$ a $G$-equivariant map. The restriction of this action onto $N$ is trivial, so it induces a group action $g_{N}\colon G_{N}\times N\backslash Y\to N\backslash Y$.
@@
Since the composition $f=f_{N}\circ q_{N}$ is continuous, we get that the map $f_{N}$ is continuous by part $(a)$ of Lemma 4.2. By part $(c)$ of Lemma 4.2 the topology on $G_{N}\times N\backslash Y$ is the quotient topology with respect to $r_{N}\times q_{N}$, so $g_{N}$ is continuous if $g_{N}\circ(r_{N}\times q_{N})$ is by part $(a)$ of Lemma 4.2. However, the composition $q_{N}\circ g$ is continuous, so by the commutativity of the following diagram:
@@
the group action $g_{N}$ is continuous. Clearly the action $g_{N}$ is free, the map $f_{N}$ is surjective, and the pre-image of each $x\in X$ is a $G_{N}$-orbit. Finally, $N\backslash Y$ is profinite by Proposition 4.5 since the action of $N$ on $Y$ is free and $N$ is compact. ∎
Definition 4.12.
Let $M,N$ be a pair of closed, normal subgroups of $G$ such that $M\subseteq N$. There is a unique map $f_{M,N}\colon M\backslash Y\to N\backslash Y$ such that $q_{N}\colon Y\to N\backslash Y$ is the composition of the quotient map $q_{M}\colon Y\to M\backslash Y$ and $f_{M,N}$. Let $\mathcal{S}$ denote the set whose elements are ordered pairs $(N,s)$, where $N$ is a closed, normal subgroup of $G$ and $s$ is a section of $f_{N}\colon N\backslash Y\to X$. Let $\geq$ denote the binary relation on $\mathcal{S}$ such that $(M,r)\geq(N,s)$ if and only if $M\subseteq N$, and $f_{M,N}\circ r=s$. Since for every triple $L\subseteq M\subseteq N$ of closed, normal subgroups of $G$ we have $f_{M,N}\circ f_{L,M}=f_{L,N}$, we get that $\geq$ is a partial ordering on $\mathcal{S}$.
@@
The poset $\mathcal{S}$ has a maximal element.
@@
For $N=G$ the map $f_{N}$ is a bijection from a compact space onto a Hausdorff topological space, so it is a homeomorphism. Therefore its inverse is a section, and hence $\mathcal{S}$ is not empty. So by Zorn’s lemma we only need to show that every chain $\mathcal{C}\subseteq\mathcal{S}$ has a maximal element. Set $C=\bigcap_{(N,s)\in\mathcal{C}}N$; since it is the intersection of closed, normal subgroups, it is also a closed, normal subgroup of $G$.
For every $(N,s)\in\mathcal{C}$, let $\Gamma(N,s)\subseteq C\backslash Y$ denote $f_{C,N}^{-1}(s(X))$, the pre-image of the section $s\colon X\to N\backslash Y$ with respect to $f_{C,N}\colon C\backslash Y\to N\backslash Y$. Since $X$ is compact and $N\backslash Y$ is Hausdorff the image $s(X)$ is closed; as $f_{C,N}$ is continuous the pre-image $\Gamma(N,s)$ is also closed. Therefore their intersection $\Gamma\subseteq C\backslash Y$ is closed, and hence compact. So it will be sufficient to show that the restriction $f_{C}|_{\Gamma}\colon\Gamma\to X$, which is continuous, is also bijective, as $X$ is Hausdorff.
Fix an $x\in X$; then for every $(N,s)\in\mathcal{C}$ the intersection $\Gamma(N,s)\cap f_{C}^{-1}(x)$ is a non-empty closed subset, and these form a descending chain with respect to inclusion. Hence by the compactness of $C\backslash Y$ their intersection $\Gamma\cap f_{C}^{-1}(x)$ is non-empty. Therefore $f_{C}|_{\Gamma}\colon\Gamma\to X$ is surjective. Now assume that we have two different $y,z\in\Gamma\cap f_{C}^{-1}(x)$; then there is a unique $1\neq\gamma\in C\backslash G$ such that $z=\gamma y$. Since $C$ is the intersection $\bigcap_{(N,s)\in\mathcal{C}}N$, there is an $(N,s)\in\mathcal{C}$ such that the image of $\gamma$ under the quotient homomorphism $C\backslash G\to N\backslash G$ is not $1$. Then the images $f_{C,N}(y)$ and $f_{C,N}(\gamma y)$ are still different, since $f_{C,N}$ is $G$-equivariant. However, both $f_{C,N}(y)$ and $f_{C,N}(\gamma y)$ lie in the intersection of $s(X)$ and $f_{N}^{-1}(x)$, which consists of the single point $s(x)$: a contradiction. Therefore $f_{C}|_{\Gamma}\colon\Gamma\to X$ is injective, too. ∎
Now let $(N,s)$ be a maximal element of $\mathcal{S}$. If we have $N=\{1\}$, then the theorem holds. So let us assume that this is not the case, and pick a non-zero $\gamma\in N$. Then there is an open normal subgroup $P\subset G$ such that $\gamma\not\in P$. Let $M=N\cap P$. Since $M$ is the intersection of closed normal subgroups, as every open subgroup is closed in a profinite group, it is such a group, and $\Gamma=M\backslash N$ is finite. Let $\Delta\subseteq M\backslash Y$ denote $f_{M,N}^{-1}(s(X))$, the pre-image of the section $s\colon X\to N\backslash Y$ with respect to $f_{M,N}\colon M\backslash Y\to N\backslash Y$. The action $g_{M}$ restricted to $\Gamma$ leaves $\Delta$ invariant.
@@
Since $f_{M}|_{\Delta}$ is the restriction of a continuous map, it is continuous, and for similar reasons $g_{M}|_{\Gamma\times\Delta}$ is continuous and $\Gamma$-equivariant, too. Since $X$ is compact and $N\backslash Y$ is Hausdorff the image $s(X)$ is closed; as $f_{M,N}$ is continuous the pre-image $\Delta$ is also closed. Therefore it is a profinite space, as $M\backslash Y$ is profinite. Finally, for every $x\in X$ the fibre $f_{M}|_{\Delta}^{-1}(x)$ is bijective $\Gamma$-equivariantly to $\Gamma$. ∎
By Lemma 4.14 and Proposition 4.9 there is a section $r\colon X\to\Delta$. Let $r$ also denote the composition of this map with the inclusion map $\Delta\to M\backslash Y$ by slight abuse of notation. Then $(M,r)\in\mathcal{S}$ such that $(M,r)\geq(N,s)$, but $(M,r)\neq(N,s)$. This contradicts the maximality of $(N,s)$, so $N=\{1\}$. This completes the proof of Theorem 4.7. ∎

### v3

@@ 4. Profinite principal $G$-bundles
[…]
@@ 1. $(a)$
    Let $A$ be a topological space and let $m\colon A\to B$ be a map. Let $B$ be equipped with the quotient topology with respect to $m$. Then a map $h\colon B\to C$ of topological spaces is continuous if and only if the composition $h\circ m\colon A\to C$ is continuous.
@@ 2. $(b)$
    Let $A$ be a topological space equipped with a left action of a group $H$. Then the quotient map $\AA\to H\backslash A$ is open, that is, it maps open sets to open sets.
@@ 3. $(c)$
[…]
@@
is continuous with respect to the product topologies, the preimage of any open subset is open. Therefore we only need to show that if $(q_{A}\times q_{B})^{-1}(U)\subseteq A\times B$ for a subset $U\subseteq H\backslash A\times H\backslash B$ is open, then $U$ is open in the product topology. In this case, $(q_{A}\times q_{B})^{-1}(U)$ is the union of sets of the form $V\times W$, where $V\subseteq A$ and $W\subseteq B$ are open. By part $(b)$ the set $(q_{A}\times q_{B})(V\times W)=q_{A}(V)\times q_{B}(W)$ is open in the product topology. Since $U$ is the union of such subsets, it is open, and $(c)$ follows. ∎
@@
Let $X$ be a profinite space and let $\mathcal{U}$ be an open covering of $X$. Then there is a finite open covering $\mathcal{V}$ of $X$ consisting of pairwise disjoint open and closed subsets which is subordinate to $\mathcal{U}$, i.e., each $V\in\mathcal{V}$ is contained in some $U\in\mathcal{U}$.
@@
By $(i)\Rightarrow(iv)$ of Theorem 3.2 and Lemma 3.7, the open and closed subsets of $X$ form a subbasis of the topology on $X$. Therefore there is an open covering $\mathcal{W}$ of $X$ consisting of open and closed subsets which is subordinate to $\mathcal{U}$. Since $X$ is compact, we may assume without the loss of generality that $\mathcal{W}$ is finite. Let $R$ be the subring of $\mathbf{B}(X)$ generated by the characteristic functions of the elements of $\mathcal{W}$. Since $R$ is finitely generated, it is finite by Corollary 2.12.
Let $\mathcal{V}$ be the collection of open and closed subsets of $X$ whose characteristic functions are the atoms of $R$. By Theorem 2.9 , we can identify them as the supports of the atoms of $R$. We claim that $\mathcal{V}$ satisfies the required properties. Because $R$ is finite, the set $\mathcal{V}$ is also finite. Since distinct atoms of $R$ are orthogonal by part $(iii)$ of Lemma 2.16, their support is pair-wise disjoint by part $(i)$ of Lemma 2.16. Let $e_{1},e_{2},\ldots,e_{n}$ be the atoms of $R$. By Corollary 2.17, we then have $1=e_{1}+\ldots+e_{n}$, so the union of their supports is $X$. Also the elements of $\mathcal{V}$ are open, since they are the supports of elements of $\mathbf{B}(X)$.
Now let $e$ be an atom of $R$ and pick an $x$ in its support $V$. Since $\mathcal{W}$ is a covering, there is a $U\in\mathcal{W}$ such that $x\in U$. The characteristic function $f$ of $U$ is in $R$, so it is the sum of distinct atoms of $R$. Therefore, $ef$ is either 0 or $e$ by the orthogonality of atoms. In the first case, the intersection of the supports $U$ and $V$ is empty, but both contain $x$, a contradiction. Hence $ef=e$, so $U$ contains $V$. Hence $\mathcal{V}$ is subordinate to $\mathcal{W}$ and since $\mathcal{W}$ is subordinate to $\mathcal{U}$, $\mathcal{V}$ is subordinate to $\mathcal{U}$. ∎
Recall that a section of a continuous map of topological spaces $f\colon Y\to X$ is a continuous map $s\colon X\to Y$ such that $f\circ s$ is the identity map of $X$.
@@
By assumption, there is an open cover $\mathcal{U}$ of $X$ such that $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$ has a section for each $U\in\mathcal{U}$. By Lemma 4.3 there is a finite open covering $\mathcal{V}$ of $X$ consisting of pairwise disjoint open and closed subsets which is subordinate to $\mathcal{U}$. Let $V\in\mathcal{V}$ and pick a $U\in\mathcal{U}$ which contains $V$. By assumption there is a section of $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$; its restriction to $V$ is a section $s_{V}\colon V\to f^{-1}(V)$ of $f|_{f^{-1}(V)}\colon f^{-1}(V)\to V$. Since the elements of $\mathcal{V}$ form an open and pairwise disjoint covering of $X$, the union $\bigcup_{V\in\mathcal{V}}s_{V}$ of these sections is a section of $f\colon Y\to X$. ∎
@@
[…]
@@
For the sake of simple notation, in the sequel we will denote any left action of any group on any set by multiplication on the left, if this does not lead to confusion. Since the action $g$ is continuous and free, for every $x\in X$ the map $\gamma\mapsto\gamma x$ from $G$ onto the $G$-orbit of $x$ is continuous and injective. As $G$ is compact and $X$ is Hausdorff, we get that the image of this map is closed by [2, I §9.4, Corollary 2 on page 87]. Moreover, the map itself is a homeomorphism from $G$ onto its image. As $G$ is homeomorphic to a closed subspace of a profinite space, it is also profinite by the equivalence $(i)\iff(ii)$ of Theorem 3.2.
[…]
Set $U^{\prime}\coloneqq\bigcup_{\gamma\in G}\gamma U$. Since each $\gamma\in G$ acts as a homeomorphism, we get that $U^{\prime}$ is the union of open subsets, so it is open. It is also the continuous image of $G\times U$. Both $G,U$ are compact since the latter is closed in a profinite space. Hence their product $G\times U$ is also compact by Tychonoff’s theorem. Thus, since $X$ is Hausdorff, $U^{\prime}$ is closed by [2, I §9.4, Corollary 2 on page 87]. Also $U^{\prime}\cap D$ is empty; if it were not then $\gamma U\cap D\neq\emptyset$ for some $\gamma\in G$, but then $U\cap D=\gamma^{-1}(\gamma U\cap D)\neq\emptyset$ using that $D$ is $G$-invariant. This is a contradiction. A similar argument shows that $V^{\prime}=\bigcup_{\gamma\in G}\gamma V$ is open and closed, and $V^{\prime}\cap C$ is empty.
[…]
@@
Let $X$ be a topological space and let $G$ be a compact group. A profinite principal $G$-bundle over $X$ is a continuous map $f\colon Y\to X$ equipped with a free continuous group action $g\colon G\times Y\to Y$ such that $Y$ is profinite, the map $f$ is surjective and the preimage of each $x\in X$ is a $G$-orbit.
@@
[…]
$(ii)$ The reader might wonder if it is true that each continuous, surjective map $f\colon X\to Y$ of profinite spaces has a section. It turns out that those $Y$ which have this property for every such $f$ have a name: extremally disconnected spaces. They can be characterised by the following property: every set of open and closed subsets of $Y$ has a supremum with respect to inclusion (see the Folk Theorem of [4] on page 485). Some spaces, such as the Čech–Stone compactification of discrete spaces have this property, but there are many profinite spaces which do not.
$(iii)$ Theorem 4.7 was already stated by Morel, see the remark after Lemma 4 of [13] on page 359. However, no proof was given, just a remark that the strategy of the proof of Proposition 1 of [21] on page 4 works. This is what we will do, but for the convenience of the reader we will give a detailed argument.
@@
By Proposition 4.4 it will be sufficient to show that every $x\in X$ has an open neighbourhood $U$ such that $f|_{f^{-1}(U)}\colon f^{-1}(U)\to U$ has a section. Since $f^{-1}(x)$ is homeomorphic to $G$, it is non-empty, so there is a $y\in Y$ such that $f(y)=x$. By [2, I §8.2, Proposition 4 on page 77], every finite subset of a Hausdorff space is closed. Because $G$ is finite and $Y$ is Hausdorff by $(i)\Rightarrow(iii)$ of Theorem 3.2, this implies that every point of $Y$ is closed. Thus, by by $(i)\Rightarrow(iv)$ of Theorem 3.2 and Lemma 3.7, there exist pairwise disjoint open and closed subsets $U_{\gamma}\subset Y$ for all $\gamma\in G$ such that $U_{\gamma}$ contains $\gamma y$ for each $\gamma$.
Set $V\coloneqq\bigcap_{\gamma\in G}\gamma^{-1}U_{\gamma}$. It is the intersection of finitely many open and closed subsets, so it is also open and closed. For every $\gamma\in G$ we have $\gamma V\subseteq\gamma(\gamma^{-1}U_{\gamma})=U_{\gamma}$, so $V\cap\gamma V=\emptyset$ for every $\gamma\neq 1$ in $G$. Therefore the restriction of $f$ to $V$ is injective. Since $V$ is closed, it is compact. Hence, since $X$ is Hausdorff by $(i)\Rightarrow(iii)$ of Theorem 3.2, $f(V)$ is closed in $X$ by [2, I §9.4, Corollary 2 on page 87]. As a closed subspace of the Hausdorff space $X$, the image $f(V)$ is Hausdorff as well. Since $V$ is compact, $f(V)$ is Hausdorff and $f|_{V}$ is a continuous bijection, $f|_{V}$ is a homeomorphism onto its image by [2, I §9.4, Corollary 2 on page 87]. Thus, $f|_{V}$ has a continuous inverse $f(V)\to V$. Since $y=\gamma^{-1}(\gamma y)\in\gamma^{-1}U_{\gamma}$, we get that $y\in\gamma^{-1}U_{\gamma}$ for every $\gamma\in G$, so $y\in V$, and hence $x\in f(V)$.
Therefore it will be sufficient to show that $f(V)$ is open in $X$. Set $Z\coloneqq\bigcup_{\gamma\in G}\gamma V$. It is clearly $G$-invariant, and since it is the union of open subsets, it is also open. Therefore its complement W=YZW=Y-Z in $Y$ is $G$-invariant and closed. Moreover, $f(V)=f(Z)$, as $f$ is $G$-equivariant, i.e., $f(\gamma y)=f(y)$ for each $\gamma\in G$ and $y\in Y$, and $f$ is surjective, so the complement of $f(V)$ in $X$ is $f(W)$. Since $W$ is closed, it is compact, and $X$ is Hausdorff, so $f(W)$ is closed by [2, I §9.4, Corollary 2 on page 87]. Therefore its complement $f(V)$ is open. ∎
Notation 4.10.
Let $N$ be a closed normal subgroup of $G$ and let $\GG_{N}=N\backslash G$ denote the quotient: it is a profinite group. Let $r_{N}\colon G\to G_{N}$ be the quotient map, which is a continuous group homomorphism. There is a unique map $f_{N}\colon N\backslash Y\to X$ such that $f\colon Y\to X$ is the composition of the quotient map $q_{N}\colon Y\to N\backslash Y$ and $f_{N}$. There is a unique group action $G\times N\backslash Y\to N\backslash Y$ which makes $f_{N}$ a $G$-equivariant map. The restriction of this action onto $N$ is trivial, so it induces a group action $g_{N}\colon G_{N}\times N\backslash Y\to N\backslash Y$.
@@
Since the composition $f=f_{N}\circ q_{N}$ is continuous, we get that the map $f_{N}$ is continuous by part $(a)$ of Lemma 4.2. By part $(c)$ of Lemma 4.2 the topology on $G_{N}\times N\backslash Y$ is the quotient topology with respect to $r_{N}\times q_{N}$, so $g_{N}$ is continuous if $g_{N}\circ(r_{N}\times q_{N})$ is continuous by part $(a)$ of Lemma 4.2. However, the composition $q_{N}\circ g$ is continuous, so by the commutativity of the following diagram:
@@
[…]
Notation 4.12.
Let $M,N$ be a pair of closed normal subgroups of $G$ such that $M\subseteq N$. There is a unique map $f_{M,N}\colon M\backslash Y\to N\backslash Y$ such that $q_{N}\colon Y\to N\backslash Y$ is the composition of the quotient map $q_{M}\colon Y\to M\backslash Y$ and $f_{M,N}$. Let $\mathcal{S}$ denote the set whose elements are ordered pairs $(N,s)$, where $N$ is a closed normal subgroup of $G$ and $s$ is a section of $f_{N}\colon N\backslash Y\to X$. Let $\geq$ denote the binary relation on $\mathcal{S}$ such that $(M,r)\geq(N,s)$ if and only if $M\subseteq N$, and $f_{M,N}\circ r=s$. Since for every triple $L\subseteq M\subseteq N$ of closed normal subgroups of $G$ we have $f_{M,N}\circ f_{L,M}=f_{L,N}$, we get that $\geq$ is a partial ordering on $\mathcal{S}$.
@@
The partially ordered set $\mathcal{S}$ has a maximal element.
@@
For $N=G$, the map $f_{N}$ is a bijection from a compact space onto a Hausdorff topological space, so it is a homeomorphism by part $(i)$ of Remark 4.8. Therefore its inverse is a section, and hence $\mathcal{S}$ is not empty. So by Zorn’s lemma we only need to show that every chain $\mathcal{C}\subseteq\mathcal{S}$ has a maximal element. Set $C\coloneqq\bigcap_{(N,s)\in\mathcal{C}}N$. Since $C$ is the intersection of closed normal subgroups, it is also a closed normal subgroup of $G$.
[…]
${C\backslash Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{C,N}}$$\scriptstyle{f_{C}}$${N\backslash Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{N}}$${X}$
[…]
${M\backslash Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{M,N}}$$\scriptstyle{f_{M}}$${N\backslash Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{N}}$${X}$
[…]
@@
[…]
Now we can finish the proof of Theorem 4.7. By Lemma 4.14 and Proposition 4.9, there is a section $r\colon X\to\Delta$. Let $r$ also denote the composition of this map with the inclusion map $\Y\Delta\to M\backslash$ Y by slight abuse of notation. Then $(M,r)\in\mathcal{S}$ and we now show that $(M,r)\geq(N,s)$. For every $x\in X$, we have $f_{N}(f_{M,N}(r(x)))=x$ since $f_{N}\circ f_{M,N}=f_{M}$ and $r$ is a section of $f_{M}$. But, by definition of $\Delta=f_{M,N}^{-1}(s(X))$, $f_{M,N}(r(x))$ is an element in $s(X)$, i.e., there is an $x^{\prime}\in X$ such that $s(x^{\prime})=f_{M,N}(r(x))$. If $x^{\prime}\neq x$, then $f_{N}(s(x^{\prime}))=x^{\prime}$ since $s$ is a section of $f_{N}$. Hence we must have $x^{\prime}=x$ and $s(x)=f_{M,N}(r(x))$, i.e., $f_{M,N}\circ r=s$. This shows $(M,r)\neq(N,s)$ in $\mathcal{S}$. However, $M$ is a proper subgroup of $N$ and hence $(M,r)\neq(N,s)$. This contradicts the maximality of $(N,s)$, so $N=\{1\}$. This completes the proof of Theorem 4.7. ∎